OpenStudy (anonymous):

Solve for X 3(2x - 2) = 2(x + 1)

OpenStudy (anonymous):

Do you know how to distribute?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

google how to distribute. getting help is okay but you should understand how i got that answer

OpenStudy (anonymous):

Thanks Again I really appreciate emily9102

OpenStudy (anonymous):

this is the solution 3(2x-2)=2(x+1) 6x-6=2x+2 collect like terms 6x-2x=2+6 4x=8 divide both sides by 4 4x/4=8/4 x=2

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

@emily9102 It's really bad to give the answer without giving any proof. I can easily say that x=3

OpenStudy (anonymous):

I know @Romero i usually ask if they need work shown. :)

OpenStudy (anonymous):

You don't give answers at all. You never learn this way. The only time you do this is if the asker has done some work or is trying to check their answer.

OpenStudy (anonymous):

And even then giving answers is not that helpful.

OpenStudy (anonymous):

@romero just wanted to double check she was more than helpful

OpenStudy (anonymous):

Okay @Romero everyone has their own way of helping. So lets just do it our own ways? Sounds good. I understand how frustrating it is not to understand a topic. So i try to help when i can. I mostly show work, but if im lazy i type answer and ask if they need work shown. Most math teachers dont accept just answers.

OpenStudy (anonymous):

dont worry emily9102 your a really good helper romero talks more than what he does

OpenStudy (anonymous):

@emily9102 That doesn't make sense....if you keep this up early in your math career you will fail eventually.

OpenStudy (anonymous):

@Romero i'm a freshman in college, if you look at my page you will see i help more than i ask.

OpenStudy (anonymous):

@emily9102 It really doesnt matter if your a freshman...your method doesn't help

OpenStudy (nastech):

x=2